OpenStudy (anonymous):

Factor the following polynomial completely. -x2y2 + x4 + 9y2 - 9x2

OpenStudy (cwrw238):

= x^4 - 9 x^2 - x^2y^2 + 9y^2 = x^2(x^2 - 9) - y^2(x^2 - 9) = (x^2 - 9)(x^2 - y^2) = (x-3)(x+3)(x-y)(x+y)